OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( -2, -5), radius = 1

OpenStudy (lalaly):

The standard form of equation of a circle is\[(x-h)^2+(y-k)^2=r^2\]where (h,k) is the center and r is the radius,, Can u figure out ur questoin now?

OpenStudy (anonymous):

Can you write the answer, please?

OpenStudy (anonymous):

\[(x+2)^{2}+(y+5)^{2}=1\]

OpenStudy (anonymous):

Thanks a lot!!